OpenStudy (anonymous):

Determine the equation of the spheres S1 and S2 of radius 2sqrt(3) that are tangent to the plane pi:x-y+z=1 at the point P=(0,0,1)

OpenStudy (anonymous):

I just solved it :D sorry =X

OpenStudy (anonymous):

but u can show urs too :D

OpenStudy (anonymous):

First the vectors : \[\Large \vec n=\begin{pmatrix}1\\-1\\1\end{pmatrix}\] and : \[\Large \vec m=-\vec n=\begin{pmatrix}-1\\1\\-1\end{pmatrix}\] are two normal vectors of the plan ! Let \[\Large A(x_A,y_A,z_A)\] The center of the sphere S1 So we can say that : \[\Large \overrightarrow{PA}=a\overrightarrow n~~~~\text{where } a>0\] To find a we have : \[\Large\|a\overrightarrow n\|=\|\overrightarrow{PA}\|=PA=R=2\sqrt3 \] So : \[\Large a=\frac{2\sqrt3}{\|\vec n\|}=\frac{2\sqrt3}{\sqrt{3}}=2\] hence : \[\Large \overrightarrow{PA}=2\overrightarrow n \] So : \[\Large \begin{cases}x_A-0&=2\\y_A-0&=-2\\z_A-1&=2\end{cases}\] So : \[\Large x_A=2~~y_A=-2~~z_A=3\] So : \[\Large A(2,-2,3).\] And then, the equation of S1is : \[\Large S_1 : ~~(x-2)^2+(y+2)^2+(z-3)^2=(2\sqrt3)^2=12\]

OpenStudy (anonymous):

great, i solved it in another way but this way is very neat too!!! congratz!

OpenStudy (anonymous):

To find the equation of s2, we can do the same but insread of using the vector n we use the vector m like this : \[\Large\text{Let }B(x_B,y_B,z_B)\text{ the center of } S_2\] We can say that : \[\Large \overrightarrow{PB}=b\overrightarrow m\] And we complete the solution as above !

OpenStudy (anonymous):

yep thats perfect thx a lot!